Boswell Company manufactures two products, Regular and Supreme. Boswell’s overhead costs consist of machining, $1000000; and assembling, $500000. Information on the two products is: Regular Supreme Direct labor hours 10000 15000 Machine hours 10000 30000 Number of parts 90000 160000 Overhead applied to Regular using traditional costing using direct labor hours is

(A) $600000.

(B) $1210000.

(C) $290000.

(D) $900000.
Answer #1

Overhead rate = estimated overhead/estimated labor hour = 1500000/25000 = 60 per DLH

Overhead applied to regular = 10000*60 = 600000

So answer is a) $600000